OpenStudy (anonymous):

If x is an acute angle of a right trangle and secx=7/5, use SOHCAHTOA to find (3sinx+4cosx)/(tanx-cotx). simplify completely, rationalizing all denominators. (*I mostly need help with simplifying the expression and rationalizing).

OpenStudy (amistre64):

there is no sohcahtoa in math ....

OpenStudy (amistre64):

sec x=7/5 cos x = 5/7 sin x = sqrt(1-(25/49)) tan x = sinx/cosx

OpenStudy (amistre64):

cot = 1/tan

OpenStudy (anonymous):

so far i have secx=7/5, therefore cosx= 5/7. sinx=2radical6/7, csc=7radical6/12. and tanx=2radical6 over 5, and cot=5radical6/12

OpenStudy (amistre64):

those look good

OpenStudy (anonymous):

plugging it into the equation is where i'm stuck. my issue is rationalizing all the radicals.

OpenStudy (amistre64):

tanx=2sqrt(6)/5 cot=5/2sqrt(6) ... yeah, same thing

OpenStudy (anonymous):

 5^2 + n^2 = 7^2 n^2 = 49 - 25 n = sqrt(24) sinx = 2sqrt(6)/7 cosx = 5/7 tanx = 2sqrt(6)/5